OpenStudy (anonymous):

what must you add to make x^2+18x=-13 a complete square?

OpenStudy (anonymous):

by adding 36?

OpenStudy (anonymous):

68

OpenStudy (anonymous):

by adding 68

OpenStudy (anonymous):

wouldnt it be 81

OpenStudy (anonymous):

it becomes (x+9)^2

OpenStudy (anonymous):

x^2+18x+13=0 (x-9)^2+13-9^2 =0 hence it should be 9^2 which is 81